# A Framework for Fractional Matrix Programming Problems with Applications in FBL MU-MIMO

### v1

@@ I Introduction
The authors of [10] proposed a quadratic transform for solving continuous fractional programming (FP) maximization problems. The algorithm proposed in [10] solves a wider range of OPs than classical FP solvers, such as Dinkelbach-based algorithms. The authors of [10] developed both power allocation and vector beamforming designs for various wireless systems, such as point-to-point communications and broadcast channels (BCs). The quadratic transform proposed in [10] introduces a new optimization variable for each FF for converting the FF to a non-fractional format. Then, alternating optimization (AO) is employed to separate the optimization of the powers (or beamforming vectors) and the optimization of the variables introduced by the transformation.
In this paper, our goal is to propose a novel unified framework for solving practical FMP OPs arising in wireless communication system design, with an emphasis on low-latency applications in MU-MIMO systems. To this end, we solve a generic minimization and a maximization problem. In these OPs, the optimization variables are matrices, in contrast to the framework in [10], which focuses on optimizing scalar and vector variables. Additionally, the authors of [10] developed a solver suitable only for maximization problems, while we propose frameworks for both maximization and minimization FMP OPs. Moreover, the OF or CFs of these FMP OPs can be arbitrary continuous functions of FFs. Furthermore, the numerator and denominator of each FF can be any arbitrary continuous function.
@@
In this paper, we first solve the generic OPs and then provide a few practical examples for each generic FMP OP, emphasizing the performance metrics that are not studied in [18, 19, 10], as illustrated in Table II. These metrics include sum delay, weighted sum EE, geometric mean of delays, geometric mean of EEs, and SEE tradeoff metrics. Additionally, we consider the MSE of MU-MIMO systems, when there are multiple data streams per user. These performance metrics are essential for enhancing the SE, EE, reliability, and latency. Furthermore, we study a broader range of MU-MIMO systems compared to [10].
@@
The paper is organized as follows. Section II presents the main results of this treatise. More specifically, Section II formulates the pair of generic FMP OPs we tackle in this paper and proposes solutions for each. Section III provides four practical examples of FMP minimization problems that our framework can solve, but Dinkelbach-based algorithms cannot. Section IV solves four commonly used FMP maximization problems that are not solvable by Dinkelbach-based algorithms. Section V shows how the framework can be extended to RIS-aided systems. Section VI extensively compares the performance and computational complexity of our FMP solver to Dinkelbach-based algorithms. Finally, Section VII concludes the paper.
@@
GDA Generalized Dinkelbach algorithm OP Optimization problem
FMP Fractional matrix programming CF Constraint function
MM Majorization minimization FF Fractional function
@@ II Problem Statement and Proposed Solutions
In this section, we formulate the OPs addressed and present our solutions. To solve these general FMP OPs, we propose iterative algorithms. Inspired by the MM framework [^1] Please refer to Appendix A for preliminaries on the MM framework., each iteration of our algorithms consists of two steps. In the first step, we reformulate the FMP problem into a simpler preferably non-FMP problem, referred to as a surrogate problem. In the second step, we solve the surrogate problem to obtain an updated point. This procedure is iterated upon fulfilling convergence criteria. Our minimization and maximization problems are discussed in separate subsections below.
@@ II-A2 General case
Note that when $f_{mi}$ is non-convex, and/or $g_{mi}$ is non-concave, we employ the following surrogate function
@@
where $\tilde{f}_{mi}(\{{\bf X}\})$ and $\tilde{g}_{mi}(\{{\bf X}\})$ fulfill the conditions in Lemma 1. Indeed, we majorize $h_{mi}(\{{\bf X}\})$ to reformulate (1) as an easier surrogate OP, as shown in Fig. 1a. Algorithm I summarizes our solution for (1) in the general case.
@@ II-B2 General case
In the framework proposed to solve (6) for the general case, we first minorize $h_{mi}(\{{\bf X}\})$ and then solve the corresponding surrogate OP in (9) to compute an SP of (6), as shown in Fig. 1b. Algorithm II summarizes our solution for (6) in the general case. Note that when $f_{mi}(\{{\bf X}\})$ is a quadratic convex function, we can directly use the inequality in (74). In this case, there is no need to use the auxiliary variable $t_{mi}$. We provide a practical example of maximizing FFs with a quadratic numerator in Section IV-D.
@@ III-A2 Rate and EE expressions
where $v_{k}$ is the channel dispersion of user $k$, ${\bf S}_{k}={\bf H}_{k}{\bf W}_{k}({\bf H}_{k}{\bf W}_{k})^{H}$ is the covariance matrix of the desired signal at user ${k}$, and ${\bf D}_{k}$ is the covariance matrix of the interfering signals plus noise, given by
@@ III-C Geometric Mean of Delays
Upon employing the inequality in Lemma 9, we have
@@ III-D Minimizing Maximum or Sum Mean Square Error
$\xi_{k}=\text{\em Tr}\left({\bf I}-{\bf D}_{k}^{-1}{\bf S}_{k}\right)\leq\tilde{\xi}_{k}=\text{\em Tr}\left({\bf I}+\tilde{\bf D}^{-1}_{k}\tilde{\bf S}_{k}\tilde{\bf D}^{-1}_{k}{\bf D}_{k}\right)\\ -2\mathfrak{R}\left\{\text{\em Tr}\left(\tilde{\bf D}^{-1}_{k}{\bf H}_{k}{\bf W}^{(z)}_{k}{\bf W}^{H}_{k}{\bf H}^{H}_{k}\right)\right\},$ (28)
@@ V Extension to RIS-aided Systems
In RIS-aided systems, we can optimize the system performance not only through the beamforming matrices $\{{\bf W}\}$ at the BS, but also through the channels by modifying the RIS coefficients. In this section, we extend our solutions in Section III and Section IV to MU-MIMO RIS-aided BCs relying on FBL coding. To this end, we employ an alternating optimization (AO) approach. More particularly, we separate the optimization of $\{{\bf W}\}$ and RIS coefficients ${\bf\Theta}$ at each iteration. Indeed, we first optimize over $\{{\bf W}\}$, when ${\bf\Theta}$ is kept fixed at ${\bf\Theta}^{(z-1)}$. Then we alternate and optimize over ${\bf\Theta}$, when $\{{\bf W}\}$ is kept fixed at $\{{\bf W}^{(z-1)}\}$. Optimization of $\{{\bf W}\}$ is similar to the solutions in Section III and Section IV. Thus, we only consider the optimization of ${\bf\Theta}$ in this section. To this end, a concave lower bound of the rates in ${\bf\Theta}$ is needed similar to when optimizing over $\{{\bf W}\}$. A concave lower bound for $\hat{r}_{k}$ is provided in [18, Corollary 1]. To avoid redundancy, we do not restate them here.
@@
where $\theta_{mn}$ denotes the entry of the $m$-th row and $n$-th column of ${\bf\Theta}$.
@@
  • RIS: The solution of our framework for RIS-aided systems.
@@ V-A Sum Delay Minimization
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ is provided in Section III-B, when sum delay is minimized. Here, we solve the sum delay minimization for fixed $\{{\bf W}\}$, which can be written as
@@
$\min_{{\bf\Theta}\in\mathcal{D},{\bf t}}\,$ $\sum_{k}\frac{L_{k}}{t_{k}}$ s.t. $\hat{r}_{k}\geq t_{k}>0,\,\,\forall k.$ (56)
The algorithm converges to an SP of the sum delay minimization OP, since the algorithm falls into the MM framework.
In Fig. 2, we show the average sum delay versus $P$, using Monte Carlo simulations. The sum delay of the algorithms considered in Fig. 2 decreases with the power budget of the BS. Moreover, the RIS substantially reduces the delay when appropriately optimized. Interestingly, to achieve a target sum delay, the RIS-aided system requires a significantly lower transmission power, which indicates how energy-efficient RIS can be, when low-latency communication is required.
@@ V-B Spectral-energy-efficiency Tradeoff
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ is provided in Section IV-A, when the SEE tradeoff metrics are maximized. Here, we provide a solution for updating ${\bf\Theta}$, when we maximize the tradeoff metric $\zeta$ stated in (36), yielding the following OP
@@
To solve (57), we employ the concave lower bound $\hat{r}_{k}$, resulting in the following convex OP
@@
The solution of (58) gives ${\bf\Theta}^{(z)}$, and the overall algorithm converges to an SP.
@@ V-C Sum EE Maximization
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ can be found in Section IV-B, when the sum EE is maximized. Employing the concave lower bound $\hat{r}_{k}$, we can update ${\bf\Theta}$ by solving

### v2

@@ I Introduction
The authors of [10] proposed a quadratic transform for solving continuous fractional programming (FP) maximization problems. The algorithm proposed in [10] solves a wider range of OPs than classical FP solvers, such as Dinkelbach-based algorithms. The authors of [10] developed both power allocation and beamforming vector designs for various wireless systems, such as point-to-point communications and broadcast channels (BCs). The quadratic transform proposed in [10] introduces a new optimization variable for each FF to convert the FF to a non-fractional format. Then, alternating optimization (AO) is employed to separate the optimization of the powers (or beamforming vectors) and the optimization of the variables introduced by the transformation.
By contrast, we propose a novel unified framework for solving practical FMP OPs arising in wireless communication system design, with an emphasis on low-latency applications in MU-MIMO systems. To this end, we solve a generic minimization and a generic maximization problem. In these OPs, the optimization variables are matrices, in contrast to the framework in [10], which focuses on optimizing scalar and vector variables. Additionally, the authors of [10] developed a solver suitable only for maximization problems, while we propose frameworks for both maximization and minimization FMP OPs. Moreover, the OF or CFs of these FMP OPs can be arbitrary continuous functions of FFs. Furthermore, the numerator and denominator of each FF can be any arbitrary continuous function.
@@
In this paper, we first solve the considered generic OPs and then provide several practical examples for each generic FMP OP, specifically emphasizing performance metrics that are not studied in [18, 19, 10], as illustrated in Table II. These metrics include the sum delay, weighted sum EE, geometric mean of delays, geometric mean of EEs, and SEE tradeoff metrics. Additionally, we consider the MSE of MU-MIMO systems, when there are multiple data streams per user. These performance metrics are essential for enhancing the SE, EE, reliability, and latency. Furthermore, we study a broader range of MU-MIMO systems compared to [10].
@@
The paper is organized as follows. Section II presents the main results of this treatise. More specifically, Section II formulates the pair of generic FMP OPs we tackle in this paper and proposes solutions for each. Section III provides four practical examples of FMP minimization problems that our framework can solve, but Dinkelbach-based algorithms cannot. Section IV solves four commonly used FMP maximization problems that are not solvable by Dinkelbach-based algorithms. Section V shows how the framework can be extended to RIS-aided systems. Section VI extensively compares the performance and computational complexity of our FMP solver to Dinkelbach-based algorithms. Section VII explores several potential directions for extending this line of research. Finally, Section VIII concludes the paper.
@@
GDA Generalized Dinkelbach algorithm HWI Hardware impairment
FMP Fractional matrix programming OP Optimization problem
MM Majorization minimization CF Constraint function
CSI Channel state information FF Fractional function
@@ II Problem Statement and Proposed Solutions
In this section, we formulate the OPs addressed and present our solutions. To solve these general FMP OPs, we propose iterative algorithms. Inspired by the MM framework [^1] Please refer to Appendix A for preliminaries on the MM framework., each iteration of our algorithms consists of two steps. In the first step, we reformulate the FMP problem into a simpler preferably non-FMP problem, referred to as a surrogate problem. In the second step, we solve the surrogate problem to obtain an updated point. This procedure is iterated until fulfilling the convergence criteria. Our minimization and maximization problems are discussed in separate subsections below.
@@ II-A2 General case
Inspired by MM, our framework utilizes an iterative approach, consisting of two steps in each iteration: majorization and minimization. Indeed, we majorize $h_{mi}(\{{\bf X}\})$ to reformulate (1) as a simpler surrogate OP, as shown in Fig. 1a. When $f_{mi}$ is non-convex and/or $g_{mi}$ is non-concave, we employ the following surrogate function
@@
where $\tilde{f}_{mi}(\{{\bf X}\})$ and $\tilde{g}_{mi}(\{{\bf X}\})$ fulfill the conditions formulated in Lemma 1. Algorithm I summarizes our solution for (1) in the general case.
@@ II-B2 General case
Similar to the framework in Section II-A, we adopt an approach inspired by MM to solve (6). In the general case, we first minorize $h_{mi}(\{{\bf X}\})$ and then solve the corresponding surrogate OP in (9) to compute an SP of (6), as shown in Fig. 1b. Algorithm II summarizes our solution for (6) in the general case. Note that when $f_{mi}(\{{\bf X}\})$ is a quadratic convex function, we can directly use the inequality in (88). In this case, there is no need to use the auxiliary variable $t_{mi}$. We provide a practical example of maximizing FFs with a quadratic numerator in Section IV-D.
@@
II-B3 Comparison with the Framework in [10]
The authors of [10] proposed a quadratic transform to solve (6) iteratively. Upon using the framework in [10] in each iteration, (6) is transformed to
$\max_{\{{\bf X}\}\in\mathcal{X},{\bf t}}$ $\sum_{m=1}^{M_{0}}\left(2t_{m0}\sqrt{{f}_{m0}(\{{\bf X}\})}-t_{m0}^{2}{g}_{m0}(\{{\bf X}\})\right)$ (10a)
s.t. $\sum_{m=1}^{M_{i}}\!\!\left(2t_{mi}\sqrt{{f}_{mi}(\{{\bf X}\})}-t_{mi}^{2}{g}_{mi}(\{{\bf X}\})\right),\,\forall i,\!$ (10b)
$t_{mi}\geq 0,\,\forall m,i,$ (10c)
which requires a joint optimization of $\{{\bf X}\}$ and ${\bf t}$. Unfortunately, (10) is non-convex even when ${f}_{mi}(\{{\bf X}\})$ and ${g}_{mi}(\{{\bf X}\})$ are linear in $\{{\bf X}\}$. To solve (10), one should employ alternating optimization to separate the optimization of ${\bf X}$ and ${\bf t}$. That is, (10) is solved by optimizing ${\bf X}$, while ${\bf t}$ is kept fixed. Then ${\bf t}$ is updated by solving (10) when ${\bf X}$ is kept fixed. Hence, this approach requires solving a pair of OPs to update $\{{\bf X}\}$ in each iteration. By contrast, our framework requires solving only a single OP in each iteration. Additionally, in the general case where ${f}_{mi}(\{{\bf X}\})$ is non-concave, it is challenging to obtain a suitable concave lower bound for $\sqrt{{f}_{mi}(\{{\bf X}\})}$. However, this issue is addressed in our framework, since Lemma 2 requires only concave lower bounds for ${f}_{mi}(\{{\bf X}\})$.
Note that the framework in [10] is proposed for solving only maximization OPs. However, we propose two frameworks for both maximization and minimization OPs. We are unaware of any other framework capable of solving the minimization FMP OPs that our framework in Section II-A can solve. Additionally, we consider practical examples for MU-MIMO FBL systems, while [10] focuses on different network scenarios when solving concrete maximization problems.
@@ III-A2 Rate and EE expressions
where $Q^{-1}$ is the inverse Q-function for Gaussian signals, $v_{k}$ is the channel dispersion of user $k$, ${\bf S}_{k}={\bf H}_{k}{\bf W}_{k}({\bf H}_{k}{\bf W}_{k})^{H}$ is the covariance matrix of the desired signal at user ${k}$, and ${\bf D}_{k}$ is the covariance matrix of the interfering signals plus noise, given by
@@ III-C Geometric Mean of Delays
The OF of (24) is a multiplication of FFs, making (24) more complex than (21). To solve (24), we first employ the upper bound in Lemma 9, transferring (24) to the format of (1). Note that the upper bound in Lemma 9 fulfills the three conditions of MM algorithms mentioned in Lemma 5, hence ensuring convergence to an SP of (24). Upon employing the inequality in Lemma 9, we have
@@ III-D Minimizing Maximum or Sum Mean Square Error
[…]
@@ V Extension to RIS-aided Systems
In RIS-aided systems, we can optimize the system performance not only through the beamforming matrices $\{{\bf W}\}$ at the BS, but also through the channels by modifying the RIS coefficients. In this section, we extend our solutions in Section III and Section IV to MU-MIMO RIS-aided BCs relying on FBL coding. To this end, we employ an alternating optimization (AO) approach, which is a typical technique of developing resource allocation schemes for multi-user RIS-assisted systems. More particularly, we separate the optimization of $\{{\bf W}\}$ and RIS coefficients ${\bf\Theta}$ at each iteration. Indeed, we first optimize over $\{{\bf W}\}$, when ${\bf\Theta}$ is kept fixed at ${\bf\Theta}^{(z-1)}$. Then we alternate and optimize over ${\bf\Theta}$, when $\{{\bf W}\}$ is kept fixed at $\{{\bf W}^{(z-1)}\}$. Optimization of $\{{\bf W}\}$ is similar to the solutions detailed in Section III and Section IV. Thus, we only consider the optimization of ${\bf\Theta}$ in this section. To this end, a concave lower bound of the rates in ${\bf\Theta}$ is needed similar to when optimizing over $\{{\bf W}\}$. A concave lower bound for $\hat{r}_{k}$ is provided in [18, Corollary 1]. To avoid redundancy, we do not restate them here.
@@
where $\theta_{mn}$ denotes the entry of the $m$-th row and $n$-th column of ${\bf\Theta}$. Note that we also consider non-convex feasibility sets for ${\bf\Theta}$ in Section V-D.
@@
  • RIS (or RISi): The solution of our framework for RIS-aided systems with feasibility set $\mathcal{D}$ (or $\mathcal{D}_{i}$).
@@ V-A Sum Delay Minimization
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ is provided in Section III-B, when the sum delay is minimized. Indeed, solving (22) gives $\{{\bf W}^{(z)}\}$. Here, we solve the sum delay minimization for fixed $\{{\bf W}\}$, which can be written as
@@
$\min_{{\bf\Theta}\in\mathcal{D},{\bf t}}\,$ $\sum_{k}\frac{L_{k}}{t_{k}}$ s.t. $\hat{r}_{k}\geq t_{k}>0,\,\,\forall k,$ (57)
whose solution is ${\bf\Theta}^{(z)}$. The algorithm converges to an SP of the sum delay minimization OP, since the algorithm falls into the MM framework.
@@ V-B Spectral-energy-efficiency Tradeoff
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ is provided in Section IV-A, when the SEE tradeoff metrics are maximized. More specifically, $\{{\bf W}^{(z)}\}$ is obtained by solving (22). Here, we provide a solution for updating ${\bf\Theta}$, when maximizing the tradeoff metric $\zeta$ stated in (37), yielding the following OP
@@
To solve (58), we employ the concave lower bound $\hat{r}_{k}$, leading to the following convex OP
@@
The solution of (59) yields ${\bf\Theta}^{(z)}$, and the overall algorithm converges to an SP.
@@ V-C Sum EE Maximization
The optimization of $\{{\bf W}\}$ for fixed ${\bf\Theta}$ can be found in Section IV-B, when the sum EE is maximized. Particularly, $\{{\bf W}^{(z)}\}$ is calculated by solving (45). Upon employing the concave lower bound $\hat{r}_{k}$, we can compute ${\bf\Theta}^{(z)}$ by solving
@@
V-D Other Feasibility Sets for the RIS Coefficients
In this subsection, we detail how one can leverage the frameworks proposed in this paper to optimize non-convex sets of RIS coefficients. To this end, we present three additional feasibility sets for the scattering matrix of nearly passive RISs, based on the models in [27, Sec. II.B]. Then, we propose a suboptimal approach to ‘convexify’ these non-convex sets. Note that the other parts of our solutions, including calculating concave lower bounds for the rates and transforming a complicated FMP OP to a simpler surrogate OP, remain unchanged, when considering a non-convex set for the feasible values of ${\bf\Theta}$. To elaborate on this issue, we solve the surrogate OP in (56) with these non-convex sets. Additionally, note that the algorithms converge, since they generate a monotone sequence of the OFs. However, we do not make any claim on the optimality of the resulting solutions, when the feasibility set of ${\bf\Theta}$ is non-convex.
Firstly, we consider the following set satisfying the unit modulus constraint:
$\mathcal{D}_{1}=\left\{\theta_{mn}:|\theta_{mm}|=1,\theta_{mn}=0,\forall m\neq n\right\}.$ (61)
The constraint $|\theta_{mm}|=1$ can be written as the two constraints $|\theta_{mm}|^{2}\leq 1$ and $|\theta_{mm}|^{2}\geq 1$. The former is convex, but the latter is not. To convexify constraint $|\theta_{mn}|^{2}\geq 1$, we relax the constraint and employ the lower bound in Lemma 11, yielding
$|\theta_{mm}|^{2}\!\geq\!|\theta_{mm}^{(z-1)}|^{2}\!-2\mathfrak{R}\{\theta_{mm}^{(z-1)^{*}}(\theta_{mm}-\theta_{mm}^{(z-1)})\}\!\geq 1-\varsigma,\!\!$ (62)
where $\varsigma>0$. Upon using (62), the resultant surrogate OP for $\mathcal{D}_{1}$ is
[…]
$\hat{r}_{k}\geq t_{k}>0,\,\,\forall k,$ (63b)
whose solution is denoted by $\hat{{\bf\Theta}}=\text{diag}(\hat{\theta}_{11},\hat{\theta}_{22},\cdots,\hat{\theta}_{MM})$, where $M$ is the number of RIS elements, and $\text{diag}(\cdot)$ denotes a diagonal matrix. Due to the relaxation in (62), $\hat{{\bf\Theta}}$ may not be in $\mathcal{D}_{1}$. To address this issue, we project $\hat{{\bf\Theta}}$ to $\mathcal{D}_{1}$ by normalizing its diagonal elements as $\hat{\theta}_{mm}^{\text{new}}=\frac{\hat{\theta}_{mm}}{|\hat{\theta}_{mm}|}$ for all $m$. To ensure convergence, we update ${\bf\Theta}$ as
$\bm{\Theta}^{(z)}=\left\{\begin{array}[]{lcl}\hat{\bm{\Theta}}^{\text{new}}&\text{if}&\sum_{k}d_{k}\left(\left\{\mathbf{W}^{(z)}\right\},\hat{{\bf\Theta}}^{\text{new}}\right)\leq\\ &&\sum_{k}d_{k}\left(\left\{\mathbf{W}^{(z)}\right\},{\bf\Theta}^{(z-1)}\right)\\ \bm{\Theta}^{(z-1)}&&\text{Otherwise},\end{array}\right.$ (64)
where $\hat{\bm{\Theta}}^{\text{new}}=\text{diag}(\hat{\theta}_{11}^{\text{new}},\hat{\theta}_{22}^{\text{new}},\cdots,\hat{\theta}_{MM}^{\text{new}})$. The updating rule in (64) guarantees convergence because a sequence of non-increasing $\sum_{k}d_{k}$ is generated.
Another set considered is based on the model in [28], assuming that the amplitude of each RIS coefficient is a deterministic function of its phase as
$\mathcal{F}(\angle\theta_{mi})=|\theta|_{\min}+(1\!-|\theta|_{\min})\left(\!\!\frac{\sin\left(\angle\theta_{mi}-\varphi\right)+1}{2}\!\right)^{\varpi}\!\!\!,\!\!$ (65)
where $|\theta|_{\min}$, $\varpi$, and $\varphi$ are non-negative constant values. Additionally, $\angle x$ takes the phase of $x$. This set can be written as [29, Eq. (7)]
$\mathcal{D}_{2}=\left\{\theta_{mn}\!:|\theta_{mm}|=\mathcal{F}(\angle\theta_{mm}),\,\angle\theta_{mm}\in[-\pi,\pi],\right.\\ \left.\theta_{mn}=0,\forall m\neq n\right\}.$ (66)
To ‘convexify’ $\mathcal{D}_{2}$, we first relax the dependency of the amplitude of $\theta_{mm}$ on its phase, resulting in
$|\theta|_{\min}^{2}\leq|\theta_{mm}|^{2}$ $\leq 1,\,\,\,\,\forall m,$ (67)
where $|\theta_{mm}|^{2}\leq 1$ is a convex constraint, but $|\theta|_{\min}^{2}\leq|\theta_{mm}|^{2}$ is non-convex. To tackle this issue, we derive a concave lower bound for $|\theta_{mn}|^{2}$ using Lemma 11, yielding the convex constraint
$|\theta_{mm}^{(z-1)}|^{2}+2\mathfrak{R}\left(\theta_{mm}^{(z-1)}(\theta_{mm}-\theta_{mm}^{(z-1)})^{*}\right)\geq|\theta|_{\min}^{2}.$ (68)
Finally, the surrogate OP for $\mathcal{D}_{2}$ is
[…]
with the solution $\hat{{\bf\Theta}}=\text{diag}(\hat{\theta}_{11},\hat{\theta}_{22},\cdots,\hat{\theta}_{MM})$. To ensure a feasible solution, we project $\hat{{\bf\Theta}}$ into $\mathcal{D}_{2}$ by choosing $\hat{{\bf\Theta}}^{\text{new}}=\text{diag}(\hat{\theta}^{\text{new}}_{11},\hat{\theta}^{\text{new}}_{22},\cdots,\hat{\theta}^{\text{new}}_{MM})$, where $\hat{{\theta}}^{\text{new}}_{mm}=\mathcal{F}(\angle\hat{{\theta}}_{mm})$ for all $m$, where $\mathcal{F}$ is defined as in (65). Finally, ${\bf\Theta}$ is updated based on (64) to ensure convergence.
The other feasibility set considered assumes a discrete domain for the phase of the RIS coefficients, while the amplitude is kept fixed at $1$ as
$\mathcal{D}_{3}=\left\{\theta_{mn}:|\theta_{mm}|=1,\angle\theta_{mm}=\{\phi_{1},\phi_{2},\cdots,\phi_{I}\},\right.\\ \left.\theta_{mn}=0,\forall m\neq n\right\},$ (70)
where the $\phi_{i}$s for all $i$ are the only possible phase shifts that can be tuned [27]. To obtain a suboptimal solution for $\mathcal{D}_{3}$, we first relax the assumption of having a discrete domain for the phase of $\theta_{mm}$. By this relaxation, ${\bf\Theta}$ can take any value in $\mathcal{D}_{1}$. Then, we compute $\hat{{\bf\Theta}}$ by solving (63). Additionally, we project $\hat{{\bf\Theta}}$ into $\mathcal{D}_{3}$ by firstly normalizing its amplitude and then taking the closest phase to $\angle\hat{\theta}_{mm}$ in $\mathcal{D}_{3}$. The projection of $\hat{{\bf\Theta}}$ in $\mathcal{D}_{3}$ is denoted by $\hat{{\bf\Theta}}^{\text{new}}$. Finally, ${\bf\Theta}^{(z)}$ is calculated according to the rule in (64) to guarantee convergence.
In Fig. 4, we show the average sum delay versus $P$, using Monte Carlo simulations. The sum delay of the algorithms considered in Fig. 4 decreases with the power budget of the BS. Moreover, for each feasibility set, the RIS substantially reduces the delay, when appropriately optimized. Interestingly, to achieve a target sum delay, the RIS-aided system requires a significantly lower transmission power, which indicates how energy-efficient RIS can be, when low-latency communication is required. Note that $I=8$ for $\mathcal{D}_{3}$ in this example, meaning that the phase of each RIS coefficient in $\mathcal{D}_{3}$ can only take values $\frac{i\pi}{4}$ for $i=1,2,\cdots,8$.